OpenStudy (anonymous):

What is the answer to this one: Log(7)231? I think I need calculator etiquette...

OpenStudy (anonymous):

Log(7)231=Log(7)[ 33 * 7] = log(7)7+log(7)33=1+log(7)33= 1+ 1.8=2.8

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

http://www.1728.org/logrithm.htm

OpenStudy (anonymous):

Samething, isaiah.

OpenStudy (anonymous):

how do you put it in the calculator, I typed it like this: pressed log (7)231 and got 195. something!!!

OpenStudy (anonymous):

No... use this calculator.. http://www.1728.org/logrithm.htm

OpenStudy (anonymous):

See how it works?

OpenStudy (phi):

you can use this formula \[ \log_b(x) = \frac{\log_{10}(x)}{\log_{10}{b}}\]

OpenStudy (phi):

your calculator knows log base 10 You could also use ln instead of log ... you should get the same answer

OpenStudy (phi):

As you noticed, your calculator does not have a log base 7... You can work around that by using the change of base formula posted up above... replace "b" with 7 and x with 231

OpenStudy (anonymous):

http://www.1728.org/logrithm.htm has any base though, you can click other and type in te base:)